OpenStudy (anonymous):

Find the value of x when S is a local minimum, justifying that it is a minimum S=2(π-2sinx)

OpenStudy (anonymous):

dS/dx = -4 cos x = 0 at critical point cos x = 0 x = pi/2 d2s/dx^2 = -4 * - sin x = 4 sin x 4 sin pi/2 is positive ( =4) so s is a minimum at x = pi/2.

OpenStudy (anonymous):

ok with that?

OpenStudy (anonymous):

Not fully

OpenStudy (anonymous):

which part are you unhappy with?

OpenStudy (anonymous):

The d then how did pi disapear. (mostly the first part)

OpenStudy (anonymous):

oh ok - sorry to be so long replying dS/dx means the derivative of S with respect to x pi is a constant and the derivative of a constant is zero ds/dx gives the slope of the tangent to the curve. this equals 0 at a critical (turning) point. eg maximum, minimum or point of inflection.

OpenStudy (anonymous):

Oh, I get it now thank you.

OpenStudy (anonymous):

yw